24 14 2. Determine the magnitude and direction (clockwise or counter-clockwise) of the resultant moment about point 0 10 ft 00 250 b 30° 3 4 ft 6 ft 3 ft 300 1b TCM 2235 Fall 2019
Answer #1

the moment or torque about the point O due to both the forces will be clock wise....乙2299-04 1b ft (Cu) 5 = Sin (6-4) b= 30-Y-30-26.57. = 3.43.